For what real x, can we have the inequality I4x-x2-3I / √(x+1) >0


1
Expert's answer
2021-08-26T13:20:12-0400

Solution:

"\dfrac{\left|4x-x^2-3\right|}{\sqrt{x+1}}>0
\\ \Rightarrow\dfrac{\left|-x^2+4x-3\right|}{\sqrt{x+1}}>0
\\ \Rightarrow\dfrac{\left|-x^2+3x+x-3\right|}{\sqrt{x+1}}>0
\\ \Rightarrow\dfrac{\left|(-x+1)(x-3)\right|}{\sqrt{x+1}}>0"

Identifying the intervals: "-1<x<1\quad \mathrm{or}\quad \:1<x<3\quad \mathrm{or}\quad \:x>3"

Finding non-negative value for radical: "x\ge-1"

Merging overlapping intervals, the solution is "\left(-1,\:1\right)\cup \left(1,\:3\right)\cup \left(3,\:\infty \:\right)"